# Counterfactual Credit Guided Bayesian Optimization

### v1

@@ B.2 Regret Bound
$\leq\Bigl(\tfrac{B_{t}}{A_{t}}\Bigr)\bigl[\mu_{t-1}(\mathbf{x}_{t})+\beta_{t}\,\sigma_{t-1}(\mathbf{x}_{t})\bigr]$
$\quad-\bigl[\mu_{t-1}(\mathbf{x}_{t})-\beta_{t}\,\sigma_{t-1}(\mathbf{x}_{t})\bigr]$
$\leq\Bigl(\tfrac{B_{t}}{A_{t}}\Bigr)2\,\beta_{t}\,\sigma_{t-1}(\mathbf{x}_{t})=\Bigl(\tfrac{B_{t}}{A_{t}}\Bigr)\,r_{t}^{\mathrm{UCB}}.$
where we define $r_{t}^{\rm UCB}\;=\;2\,\beta_{t}\,\sigma_{t-1}(\mathbf{x}_{t})$ as the standard UCB instantaneous regret bound. Summing over $t=1,\dots,N$ gives
$\begin{split}R_{N}^{\rm ccg}&=\sum\nolimits_{t=1}^{N}r_{t}^{\rm ccg}\\ &\leq\max_{1\leq t\leq N}\!\Bigl(\tfrac{B_{t}}{A_{t}}\Bigr)\sum\nolimits_{t=1}^{N}r_{t}^{\rm UCB}\\ &=\max_{1\leq t\leq N}\!\Bigl(\tfrac{B_{t}}{A_{t}}\Bigr)\;R_{N},\end{split}$ (15)
and the result follows by the standard bound $R_{N}=O(\sqrt{N\,\gamma_{N}\,\beta_{N}})$. ∎
@@ Appendix C CCGBO with other acquisition functions
CCGBO is a plug-and-play module and can be combined with acquisition functions beyond UCB. The counterfactual credit-based weight can be simply multiplied by other acquisition functions as a coefficient. As suggested, we have evaluated it with several others, including TS in Figure 4, logEI [Ament et al., 2023] in Figure 5, and JES [Hvarfner et al., 2022a] in Figure 6. We observe that CCGBO yields similar performance gains when paired with JES and TS, just as with UCB, whereas with logEI (an EI-family variant) the improvement is negligible. We believe this difference stems from the inherently greedy nature of EI-style methods: since they already concentrate heavily on nearby “promising” regions, multiplying by credit weights does not materially change their ranking or introduce substantial additional guidance. In contrast, acquisitions with explicit or implicit exploration (e.g., UCB, TS, JES) benefit more from the extra bias toward high-contribution regions while retaining their ability to explore.
@@ Appendix D High-dimensional Tasks
However, directly comparing against algorithms specifically designed for high‑D BO would be unfair, since those methods adapt their kernels or acquisition functions to that setting. Instead, we adopted the straightforward “Vanilla Bayesian Optimization Performs Great in High Dimensions” [Hvarfner et al., 2024a] strategy: in a standard GP + acquisition framework, one simply shifts the log‐lengthscale prior upward by $\tfrac{1}{2}\log D$ (i.e. scales the lengthscale by $\sqrt{D}$) and fixes the signal variance to 1. We applied this modified GP backbone across all CCGBO and baseline algorithms to ensure a fair comparison.
In the experiments, we follow [Hvarfner et al., 2024a] to employ Embedded Test Functions to evaluate the performance of Bayesian optimization algorithms in high-dimensional spaces. These functions are constructed by embedding classical low-dimensional benchmark functions into higher-dimensional spaces. Given a $d_{eff}$-dimensional benchmark function $f:\mathbb{R}^{d_{eff}}\rightarrow\mathbb{R}$, we embed it into a $D$-dimensional space ($D\gg d_{eff}$) to construct a new objective function $f_{embed}:\mathbb{R}^{D}\rightarrow\mathbb{R}$: $f_{embed}(\mathbf{x})=f(\mathbf{x}_{1:d_{eff}})$ where $\mathbf{x}_{1:d_{eff}}$ denotes the first $d_{eff}$ components of the input vector $\mathbf{x}\in\mathbb{R}^{D}$. This means that the objective function value depends only on the first $d_{eff}$ dimensions (effective dimensions), while the remaining $D-d_{eff}$ dimensions have no effect on the function value. In this work, two classical benchmark functions are adopted as bases: (1) Levy function: $4$ effective dimensions, characterized by multiple local optima, used to evaluate global search capability. (2) Hartmann function: $6$ effective dimensions, with a complex non-convex landscape, widely used in optimization benchmarking. These functions are embedded into $25$, $100$, $300$, and $1000$-dimensional spaces for experiments. For example, Levy4_1000 denotes embedding the $4$-dimensional Levy function into a $1000$-dimensional space.

### v2

@@ B.2 Regret Bound
$=\bigl(f(\mathbf{x}^{*})-B_{\min,t}\bigr)-\bigl(f(\mathbf{x}_{t})-B_{\min,t}\bigr)$
$\leq\bigl[\mu_{t-1}(\mathbf{x}^{*})+\beta_{t}\,\sigma_{t-1}(\mathbf{x}^{*})-B_{\min,t}\bigr]-\bigl[f(\mathbf{x}_{t})-B_{\min,t}\bigr]$
$\leq\Bigl(\tfrac{B_{t}}{A_{t}}\Bigr)\bigl[\mu_{t-1}(\mathbf{x}_{t})+\beta_{t}\,\sigma_{t-1}(\mathbf{x}_{t})-B_{\min,t}\bigr]-\bigl[\mu_{t-1}(\mathbf{x}_{t})-\beta_{t}\,\sigma_{t-1}(\mathbf{x}_{t})-B_{\min,t}\bigr]$
$=\underbrace{\Bigl(\tfrac{B_{t}}{A_{t}}\Bigr)\,2\beta_{t}\,\sigma_{t-1}(\mathbf{x}_{t})}_{\text{standard term}}+\underbrace{\Bigl(\tfrac{B_{t}}{A_{t}}-1\Bigr)\bigl[\mu_{t-1}(\mathbf{x}_{t})-\beta_{t}\,\sigma_{t-1}(\mathbf{x}_{t})\bigr]}_{\text{residual term}}-\underbrace{\Bigl(\tfrac{B_{t}}{A_{t}}-1\Bigr)B_{\min,t}}_{\text{shift term}}.$
For the standard term, note that $c_{\min}>0$ and $c_{\max}\leq 1$ imply that $A_{t}$ and $B_{t}$ are uniformly bounded away from 0 and $\infty$, so there exists a constant $C_{1}>0$ such that $\frac{B_{t}}{A_{t}}\leq C_{1}$ for all $t$. Hence
$\frac{B_{t}}{A_{t}}\,2\beta_{t}\sigma_{t-1}(\mathbf{x}_{t})\;\leq\;2C_{1}\,\beta_{t}\sigma_{t-1}(\mathbf{x}_{t}).$
Summing over $t$ and using the standard GP-UCB argument of Srinivas et al. (2010), we obtain
$\sum_{t=1}^{N}\frac{B_{t}}{A_{t}}\,2\beta_{t}\sigma_{t-1}(\mathbf{x}_{t})=O\bigl(\sqrt{N\,\gamma_{N}\,\beta_{N}}\bigr).$
For the residual and shift terms, observe that $\mu_{t-1}(\mathbf{x}_{t})-\beta_{t}\sigma_{t-1}(\mathbf{x}_{t})$ is the lower confidence bound. Under Assumption 2, we have $|f(\mathbf{x})|\leq B$ for all $\mathbf{x}$, and on the confidence event
$\mu_{t-1}(\mathbf{x}_{t})-\beta_{t}\sigma_{t-1}(\mathbf{x}_{t})\;\leq\;f(\mathbf{x}_{t})\;\leq\;B.$
For the shift term, note that on the confidence event $\mu_{t-1}(\mathbf{x})+\beta_{t}\sigma_{t-1}(\mathbf{x})\geq f(\mathbf{x})\geq-B$ for all $\mathbf{x}$. Hence $B_{\min,t}\geq-B$, which implies $B_{\min,t}\leq B$.
Moreover, by construction $B_{t}\geq A_{t}$, so $\frac{B_{t}}{A_{t}}-1\geq 0$. From the explicit expressions
$A_{t}=1-\lambda+\lambda c_{\min}^{\alpha_{t}},\qquad B_{t}=1-\lambda+\lambda c_{\max}^{\alpha_{t}},\qquad\alpha_{t}=\frac{\tau}{1+t/M},$
and the fact that $c_{\min},c_{\max}\in(0,1]$, a simple application of the mean value theorem shows that there exists a constant $C_{2}>0$ such that
$\frac{B_{t}}{A_{t}}-1\;\leq\;\frac{C_{2}}{1+t/M}\quad\text{for all }t.$
Therefore, combining the residual and shift terms and using the bounds above:
$\sum_{t=1}^{N}\Bigl(\frac{B_{t}}{A_{t}}-1\Bigr)\bigl[\mu_{t-1}(\mathbf{x}_{t})-\beta_{t}\sigma_{t-1}(\mathbf{x}_{t})\bigr]-\sum_{t=1}^{N}\Bigl(\frac{B_{t}}{A_{t}}-1\Bigr)B_{\min,t}\;\leq\;\sum_{t=1}^{N}2\Bigl(\frac{B_{t}}{A_{t}}-1\Bigr)B\;\leq\;2C_{2}B\sum_{t=1}^{N}\frac{1}{1+t/M}=O(\log N).$
Combining the bounds for the standard, residual, and shift terms, we get
$R_{N}^{\mathrm{ccg}}=\sum_{t=1}^{N}r_{t}^{\mathrm{ccg}}=O\bigl(\sqrt{N\,\gamma_{N}\,\beta_{N}}\bigr)+O(\log N)=O\bigl(\sqrt{N\,\gamma_{N}\,\beta_{N}}\bigr),$
so CCGBO retains the same sublinear regret rate as GP-UCB.
@@ Appendix C CCGBO with other acquisition functions
Relation to Information-based Acquisition Functions.
Information-based acquisition functions, such as Entropy Search (ES) (Hennig and Schuler, 2012), Predictive Entropy Search (PES) (Hernández-Lobato et al., 2014), and Max-value Entropy Search (MES) (Wang and Jegelka, 2017), focus on selecting points that maximize the information gain about the location of the optimum $\mathbf{x}^{*}$. While conceptually related to our work in seeking to identify promising regions, CCGBO differs in two key aspects: (1) Our core principle is to use the function value proxy $Z_{t}$ to compute counterfactual credit for reweighting the acquisition function, rather than directly modeling the distribution of $\mathbf{x}^{*}$; information-based methods still treat all candidate points equally when computing information gain. (2) CCGBO is a generic upper-layer module that can be combined with any base acquisition function, including ES-family methods, whereas ES/PES/MES are standalone acquisition functions. We demonstrate this compatibility by combining CCGBO with Joint Entropy Search (JES) (Hvarfner et al., 2022a), an advanced member of the ES family, and observe consistent improvements (see Figure 6).
Compatibility with Different Acquisition Functions.
CCGBO is a plug-and-play module and can be combined with acquisition functions beyond UCB. The counterfactual credit-based weight can be simply multiplied by other acquisition functions as a coefficient. As suggested, we have evaluated it with several others, including TS in Figure 4, logEI (Ament et al., 2023) in Figure 5, and JES (Hvarfner et al., 2022a) in Figure 6. We observe that CCGBO yields similar performance gains when paired with JES and TS, just as with UCB, whereas with logEI (an EI-family variant) the improvement is negligible. We believe this difference stems from the inherently greedy nature of EI-style methods: since they already concentrate heavily on nearby “promising” regions, multiplying by credit weights does not materially change their ranking or introduce substantial additional guidance. In contrast, acquisitions with explicit or implicit exploration (e.g., UCB, TS, JES) benefit more from the extra bias toward high-contribution regions while retaining their ability to explore.
@@ Appendix D High-dimensional Tasks
However, directly comparing against algorithms specifically designed for high‑D BO would be unfair, since those methods adapt their kernels or acquisition functions to that setting. Instead, we adopted the straightforward “Vanilla Bayesian Optimization Performs Great in High Dimensions” (Hvarfner et al., 2024a) strategy: in a standard GP + acquisition framework, one simply shifts the log‐lengthscale prior upward by $\tfrac{1}{2}\log D$ (i.e. scales the lengthscale by $\sqrt{D}$) and fixes the signal variance to 1. We applied this modified GP backbone across all CCGBO and baseline algorithms to ensure a fair comparison.
In the experiments, we follow (Hvarfner et al., 2024a) to employ Embedded Test Functions to evaluate the performance of Bayesian optimization algorithms in high-dimensional spaces. These functions are constructed by embedding classical low-dimensional benchmark functions into higher-dimensional spaces. Given a $d_{eff}$-dimensional benchmark function $f:\mathbb{R}^{d_{eff}}\rightarrow\mathbb{R}$, we embed it into a $D$-dimensional space ($D\gg d_{eff}$) to construct a new objective function $f_{embed}:\mathbb{R}^{D}\rightarrow\mathbb{R}$: $f_{embed}(\mathbf{x})=f(\mathbf{x}_{1:d_{eff}})$ where $\mathbf{x}_{1:d_{eff}}$ denotes the first $d_{eff}$ components of the input vector $\mathbf{x}\in\mathbb{R}^{D}$. This means that the objective function value depends only on the first $d_{eff}$ dimensions (effective dimensions), while the remaining $D-d_{eff}$ dimensions have no effect on the function value. In this work, two classical benchmark functions are adopted as bases: (1) Levy function: $4$ effective dimensions, characterized by multiple local optima, used to evaluate global search capability. (2) Hartmann function: $6$ effective dimensions, with a complex non-convex landscape, widely used in optimization benchmarking. These functions are embedded into $25$, $100$, $300$, and $1000$-dimensional spaces for experiments. For example, Levy4_1000 denotes embedding the $4$-dimensional Levy function into a $1000$-dimensional space.
@@ Appendix G Ablation Study on Nearest Neighbors H
Appendix H Ablation Study on Additional Hyperparameters
In this section, we provide ablation studies on the remaining hyperparameters: the credit influence parameter $\lambda$, the UCB exploration factor $\beta_{t}$, the numerical stability constant $\varepsilon_{c}$, the minimum credit bound $r_{\min}$, and the credit decay exponent $\tau$.
Credit Influence Parameter $\lambda$.
The parameter $\lambda\in[0,1]$ controls the strength of credit influence in the acquisition function. When $\lambda=0$, CCGBO reduces to standard UCB; when $\lambda=1$, the credit weighting is fully applied. Table 2 reports the final simple regret on the Hartmann6 benchmark for different values of $\lambda$.
Table 2: Ablation of credit influence parameter $\lambda$ on simple regret (Hartmann6, mean $\pm$ std over 50 runs).
$\lambda$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
[…]
We observe that performance improves as $\lambda$ increases from small to moderate values and maintains similar performance in the range $\lambda\geq 0.5$. This indicates that CCGBO is robust to the choice of $\lambda$ over a wide range, rather than requiring careful tuning.
Comparison with GP-UCB under Varying $\beta_{t}$.
A natural question is whether simply reducing the exploration parameter $\beta_{t}$ in standard GP-UCB could achieve comparable performance to CCGBO by inducing more exploitative behavior. To investigate this hypothesis, we conducted a systematic comparison between GP-UCB with varying $\beta_{t}$ values and CCGBO with a fixed $\beta_{t}=2.576$. Table 3 reports the final simple regret on three representative benchmark problems.
Table 3: Comparison of GP-UCB with varying $\beta_{t}$ values against CCGBO on simple regret (mean $\pm$ std over 50 runs). Lower is better.
Problem $\beta=0.5$ $\beta=1.0$ $\beta=2.0$ $\beta=2.576$ $\beta=4.6$ CCGBO
[…]
The results reveal that reducing $\beta_{t}$ alone cannot replicate the performance of CCGBO. On the unimodal Rosenbrock2 function, smaller $\beta_{t}$ values yield marginal improvements but still underperform CCGBO, which achieves both lower mean regret and substantially reduced variance. On multimodal landscapes (Hartmann6 and Levy8), decreasing $\beta_{t}$ suppresses exploration globally, causing GP-UCB to converge prematurely to local optima. In contrast, CCGBO selectively amplifies exploitation only in regions with high estimated contribution while preserving standard exploration–exploitation balance elsewhere. Furthermore, the time-decaying weighting mechanism ensures that CCGBO gradually reverts to standard UCB behavior in later iterations, thereby avoiding entrapment in local optima.
Overall, these ablation studies demonstrate that CCGBO maintains robust performance across a wide range of hyperparameter settings, indicating that the reported improvements are not artifacts of careful tuning but reflect the inherent benefits of the counterfactual credit mechanism.